OpenStudy (anonymous):

The letters of the word STAR are written on individual cards. If you always choose three cards and then put them back before the next set of three cards, how many different permutations will you have? List all the different permutations in the space provided

OpenStudy (experimentx):

what do you mean by next three set of cards?

OpenStudy (anonymous):

Here are the three cards you chose: ___ / ___ / ___ How many possibilities for the first card? How many for the second card (be careful!)? How many for the third? Multiply those up, and you'll have your answer.

OpenStudy (experimentx):

if you are choosing next 3 sets of cards like you choose your first cards then.4!x4! (if you are placing them 'before' next 3 set of cards.

OpenStudy (anonymous):

It's a really badly worded question.

OpenStudy (anonymous):

im lost too lol

OpenStudy (experimentx):

be clear of the situation in the question first.

OpenStudy (anonymous):

4! x 4! , i think...

OpenStudy (chrisneedshelp):

I MEANTTT 4x4

OpenStudy (anonymous):

I think it just means "How many different ways are there to order three cards taken from four cards?" If that's the case, then my explanation above was correct.

OpenStudy (anonymous):

stilll lost.-.

OpenStudy (ganpat):

No of ways, 3 cards can be take from 4.. 4C3 = 4 and arranging those 3 cards = 3! = 3*2*1 = 6 so total ways = 4 *6 = 24.. This might be answer..